OpenStudy (anonymous):

2/1/8 + 3/3/4 perform the indicated operation.

OpenStudy (anonymous):

16+4=20 16=2/(1/8) 4=3/(3/4)

OpenStudy (anonymous):

I got 17/12 for the final answer. can this be reduced any further?

OpenStudy (anonymous):

2/(1/8) + 3/(3/4) is it like this?

OpenStudy (anonymous):

no. sorry. mixed fractions.

OpenStudy (anonymous):

sorry. i got 47/8 for the final. yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks.